OpenStudy (anonymous):

the slope of the function y= 20x + 5 is?

OpenStudy (phi):

see http://www.mathwarehouse.com/algebra/linear_equation/slope-intercept-form.php scroll down to "Formula"

OpenStudy (anonymous):

slope is 20 units

OpenStudy (anonymous):

two ways to solve this 1) the standard form of equation of a line is y=mx+b where m is the slope by comparing the standard form with the given equation we see that the slope = 20 units 2)differentiate the equation with respect to x to obtain the slope dy/dx = d/dx(20x) +d/dx(5) dy/dx = 20